(3 square root of 6)(2 square root 2): (3 square root of 6)(2 square root 2)
Answer by checkley77(3412)

(3sqrt6)(2sqrt2)=6sqrt12=6sqrt(4*3)=(6*2)sqrt3=12sqrt3=20.7846